# Prague, Czech

### v3

@@ Abstract
We propose a reinforcement learning (RL) framework for adaptive precision tuning of linear solvers, and can be extended to general algorithms. The framework is formulated as a contextual bandit problem and solved using incremental action-value estimation with a discretized state space to select optimal precision configurations for computational steps, balancing precision and computational efficiency. To verify its effectiveness, we apply the framework to iterative refinement for solving linear systems $Ax=b$. In this application, our approach dynamically chooses precisions based on calculated features from the system. In detail, a Q-table maps discretized features (e.g., approximate condition number and matrix norm)to actions (chosen precision configurations for specific steps), optimized via an $\epsilon$-greedy strategy to maximize a multi-objective reward balancing accuracy and computational cost. Empirical results demonstrate effective precision selection, reducing computational cost while maintaining accuracy comparable to double-precision baselines. The framework generalizes to diverse out-of-sample data and offers insight into utilizing RL precision selection for other numerical algorithms, advancing mixed-precision numerical methods in scientific computing. To the best of our knowledge, this is the first work on precision autotuning with RL and verified on unseen datasets.
Keywords inference reduced-precision configurations, reinforcement learning, precision tuning, linear solvers
@@ 1 Introduction
Floating-point arithmetic (8766229, ) is widely used in machine learning and scientific computing. With the rapid development of large language models (DBLP:journals/corr/abs-2005-14165, ; hu2022lora, ; touvron2023, ) as well as large-scale GPU computing (DBLP:journals/corr/abs-1710-03740, ; peng2023fp8lmtrainingfp8large, ), numerical calculations and algorithm deployment with extensive use of floating-point arithmetic are often accompanied by a substantial increase in energy use. Therefore, using reduced-precision arithmetic in algorithms is typically required in many intensive computing kernels due to its lower communication, faster arithmetic, and energy efficiency. Seven commonly-used floating-point formats are referenced in Table 1. The NVIDIA A100 achieves up to 312 TFLOPS with half precision (FP16/BF16), 156 TFLOPS with TF32, and 9.7 TFLOPS with FP64, yielding an approximate $32\times$ speedup for low-precision formats. Further, the NVIDIA H100 (NVIDIA’s Hopper architecture) achieves up to 624 TFLOPS with fp8, offering an approximate $64\times$ speedup versus FP64 (nvidia_a100, ; nvidia_fp8, ). The work nvidia_fp8 demonstrates that the FP8 formats E4M3 and E5M2 can in some cases match FP16 accuracy with reduced compute and memory costs.
@@
However, reduced-precision arithmetic often brings increased rounding errors and numerical instability, which can compromise the results of computations, especially at large scale (10.1177/1094342016652462, ). The idea behind mixed-precision algorithms is to design an algorithm that combines low and high precision arithmetic while retaining an acceptable level of accuracy, which remains a major obstacle in modern approximate computing paradigms (benkhalifa:hal-03978176, ; 7348659, ). Accordingly, many theoretical analyses on rounding error analysis and mixed-precision algorithms (doi:10.1137/17M1122918, ; doi:10.1137/17M1140819, ; doi:10.1137/20M1316822, ; doi:10.1137/17M1122918, ; abhl24, ) and precision tuning tools (6877460, ; 10.1145/3213846.3213862, ) have been developed, which aim to achieve a balance between precision loss and arithmetic speedup.
@@
Theoretical analyses of the use of multiprecision arithmetic and mixed-precision algorithms have been extensively conducted in the context of iterative refinement (doi:10.1137/20M1316822, ; abhl24, ), least squares problems (doi:10.1137/19M1298263, ; doi:10.1137/20M1316822, ), and GMRES (jang:hal-05163845, ). For example, iterative refinement can be executed in three precisions, but the convergence is only theoretically guaranteed for a limited range of problems, which can be overly restrictive in practice oktay2022multistage . As such, one can instead use to precision tuning tools to find a suitable mixed-precision approach. Existing tools typically require running the program multiple times via trial-and-error, and the scalability and general ability are not verified, so they are either limited to small or medium-scale programs or suffer from efficiency issues. An ideal case for a precision tuning tool or method is to fit the model in a limited number of training sets and learn the key features that can be used to infer mixed-precision configurations for new unseen data. These challenges need to be addressed to reduce the dependency on programmer efforts.
@@
A contextual bandit (also known as a multi-armed bandit with context) is a simplified RL problem where the agent makes decisions based on the current context (i.e., state) without considering the future contexts or rewards of its actions; in other words, the agent optimizes for one-step (immediate) reward rather than long-term cumulative reward. In this paper, we propose a general framework for applying reinforcement learning (RL) for precision selection in general numerical algorithms, formulated as a contextual bandit problem due to the independence of problem instances and the single decision per instance. This is the first work to formally define the use of RL for precision selection in any algorithm. The framework uses an epsilon-greedy strategy, controlled by a decaying epsilon determined by training episodes, to select precision levels for the prespecified algorithm steps, treating precision choices as a multi-armed bandit problem within a RL environment. The epsilon parameter balances exploration and exploitation, starting high to encourage diverse precision selections and decaying over time to favor optimal choices. We believe this framework enables at least two utility points:
@@
    Being able to inference reduced mixed-precision configurations of unseen datasets.
We apply this framework to the GMRES-based iterative refinement (GMRES-IR) method doi:10.1137/17M1122918 to verify the effectiveness and generalization ability of RL for precision selection, aiming to achieve a trade-off between the use of low precision and accuracy. The framework is particularly suited for scenarios with limited training data and is designed to generalize to new data, and can be easily implemented in an online learning routine to avoid model retraining.
The rest of the paper is structured as follows. Section 2 discusses the existing literature on precision autotuning as well as its pros and cons, and a comparison with our work. Section 3 presents our general framework for precision selection; To verify its effectiveness, we use the iterative refinement method, a well-studied mixed-precision method, as a case study detailed in Section 4. Section 5 presents the experiments and empirical results on both dense and sparse matrices under varying condition numbers and sparsity. Section 7 discusses limitations as well as future work.
@@ 2 Related Work
There exists a variety of theoretical rounding error analysis of basic mixed-precision algorithms aimed at understanding the underlying arithmetic behavior; see the survey doi:10.1177/10943420211003313 and references therein. However, for real-world problems, which can be more complex, the scientific community typically uses more general approaches such as precision tuning tools, which operate at program compile time and runtime, immediately resulting in reduced-precision programs produced automatically. In this section, we briefly review existing approaches for the analysis of mixed-precision algorithms and automated precision tuning that are closely related to our method.
@@
Theoretical analysis aims to design reduced-precision settings for specific steps and analyze their rounding error so that the reduced-precision arithmetic can be used safely under certain assumptions to retain algorithm accuracy and numerical stability. Analyses in WILKINSON1966 and 10.1145/321386.321394 demonstrate that in standard iterative refinement, residuals computed in extra precision (typically twice the working precision) can guarantee the convergence of the relative error to the level of the working precision if the matrix $A$ has a condition number safely less than the reciprocal of the unit-roundoff, $u$. More recently, doi:10.1137/17M1122918 found that it is possible to produce solutions to iterative refinement with normwise relative error of order $u$ for systems with condition numbers of order $u^{-1}$ or larger by using a GMRES-based approach called GMRES-IR (see Algorithm 2). The authors extended this work in doi:10.1137/17M1140819 to develop a general analysis of iterative refinement using up to three difference precisions, as well as a general solver. This work shows that, depending on constraints on the condition number $\kappa(A)$, the most expensive part of the computation can often be carried out entirely in low precision without harming the accuracy.
However, to obtain theoretical results such as these, a great amount of effort is needed, and the analysis often relies on knowledge about the problem, e.g., the condition number, which may not be readily available in practice. To circumvent these obstacles, one can switch to precision tuning tools that operate during compile time and runtime; see the survey 10.1145/3381039 and references therein.
@@
Another approach is to use an automated precision tuning method. Such methods, built upon the software level via static and dynamic analysis tools (benkhalifa:hal-03978176, ), can obtain suitable reduced-precision settings through trial and error. The search-based tool Precimonious (6877460, ) uses high-performance dynamic programming approaches, which achieve a potentially lower precision setting for the program via Delta Debugging of the space of possible floating-point configurations within the target accuracy. Further, HiFPTuner (10.1145/3213846.3213862, ) (High-level Floating-Point Tuner) operates at the source level to automatically tune the floating-point precision via profiling floating-point operations and estimating the numerical impact of the reduced-precision program. CRAFT (10.1177/1094342016652462, ) builds reduced-precision configurations of floating-point programs based on the Dyninst binary instrumentation framework (10.1177/109434200001400404, ), and uses an automatic search process that identifies the precision-level sensitivity of various parts of an application. Search-based tools for finding a low-precision configuration typically require a dramatic increase in computational cost as the number of variables increases, and do not scale well to large programs. To further reduce the runtime of precision tuning, Blame Analysis (Rubio2016, ) is devoted to executing a program once, analyzing floating-point instructions to determine the minimum precision required for each operand, reducing the search space of Precimonious and improving scalability. PyFloT (9355251, ) automates mixed-precision tuning for HPC applications by leveraging runtime profiling, instruction-level granularity, and temporal locality to learn numerical behavior and recurring patterns in instruction execution. FPLearner (10.1145/3597503.3623338, ) leverages a graph representation learning technique to predict the performance and accuracy of a mixed-precision program, which can then be used to accelerate dynamic precision tuning by reducing the number of program runs.
@@ 3.1 Contextual Bandit Formulation
3.2 Learning with Discretized State Space
@@
$s_{d}=\text{discretize}(s)=\left(\text{bin}(\phi_{1}(s)),\ldots,\text{bin}(\phi_{d}(s)))\right),$ (4)
@@
To handle the large action space $|\mathcal{A}|=m^{k}$, we decompose the action selection by maintaining $k$ separate Q-tables, $Q_{i}:\mathcal{S}_{d}\times\mathcal{A}_{i}\to\mathbb{R}$, for $i=1,\ldots,k$, where $Q_{i}(s_{d},a_{i})$ estimates the expected reward for selecting precision $a_{i}\in\mathcal{A}_{i}$ for the $i$-th step in state $s_{d}$. This assumes independence between precision choices, reducing the Q-table size from $|\mathcal{S}_{d}|\cdot m^{k}$ to $k\cdot|\mathcal{S}_{d}|\cdot m$.
@@
Algorithm 1 RL Framework for Precision Selection in Numerical Algorithms
1:Input: Training instances $\mathcal{P}=\{p_{1},\ldots,p_{N}\}$, numerical method $\mathcal{M}$, precision sets $\{\mathcal{A}_{i}\}_{i=1}^{k}$, reward function $r$, episodes $T$, learning rate $\alpha$, discount factor $\gamma$, exploration epsilon $\epsilon$, bin sets $\{\Phi_{j}\}_{j=1}^{d}$
2:Initialize Q-table $Q(s_{d},i,a_{i})$ for all $s_{d}\in\mathcal{S}_{d}$, $i=1,\ldots,k$, $a_{i}\in\mathcal{A}_{i}$
3:Initialize environment $\mathcal{E}$ with instances $\{p_{i}\}$
4:for episode $t=1$ to $T$ do
5:  for each instance $p_{i}$, $i=1,\ldots,N$ do
6:   Compute state in terms of data features $s_{i}\leftarrow[\phi_{1}(p_{i}),\ldots,\phi_{d}(p_{i})]$ (e.g., log condition number, log norm)
7:   Discretize state $s_{d}\leftarrow\text{discretize}(s_{i},\{\Phi_{j}\})$
8:   Choose action $a\in\mathcal{A}$ uniformly with probability $\epsilon$ otherwise use $a\leftarrow\arg\max_{a^{\prime}\in\mathcal{A}}Q(s_{d},a^{\prime})$
@@
10:   Run $\mathcal{M}$ with precision $a$ in $\mathcal{E}$, observe reward $r(s_{i},a)$, next state $s_{i}^{\prime}$, termination status
11:   Discretize next state $s_{d}^{\prime}\leftarrow\text{discretize}(s_{i}^{\prime},\{\Phi_{j}\})$
12:   Update Q-table:
$Q(s_{d},\mathbf{a})\leftarrow=Q(s_{d},\mathbf{a})+\alpha\Big(R(s_{d},\mathbf{a})-Q(s_{d},\mathbf{a})\Big),$
13:   if termination status is true then
14:     Reset $\mathcal{E}$ to next instance
15:   end if
16:  end for
17:  Update epsilon $\epsilon$
18:end for
19:Inference: For new instance $p$, compute $s=[\phi_{1}(p),\ldots,\phi_{d}(p)]$, discretize $s_{d}=\text{discretize}(s,\{\Phi_{j}\})$, select $a_{i}=\arg\max_{a\in\mathcal{A}_{i}}Q(s_{d},i,a)$, apply $\mathcal{M}$ with precision $\mathbf{a}$
Our RL learning framework for precision selection can be extended to general algorithms by properly remodeling the state and action space. The pseudocode the linear solve is detailed in Algorithm 1. The framework aims to optimize precision usage for the $k$ computational steps of a mixed-precision algorithm applied to linear systems $p=(A,b)\in\mathcal{P}$. It iterates over a training set of $N$ problem instances for $T$ episodes, selecting actions according to the estimated Q-values. For each problem instance, it begins by computing a context $s\in\mathbb{R}^{d+n}$ from matrix properties, discretizing it into $s_{d}\in\mathcal{S}_{d}$ using bins $\Phi_{j}$. Actions $\mathbf{a}=(a_{1},\ldots,a_{k})\in\mathcal{A}$ are selected via a softmax policy $\pi(a_{i}|s_{d})\propto\exp(\text{clip}(Q_{i}(s_{d},a_{i}),-Q_{\text{max}},Q_{\text{max}})/\tau)$, where $k$ Q-tables $Q_{i}$ guide precision choices for each step. The solver $\mathcal{M}$ is applied with $\mathbf{a}$, yielding a reward $r(s,\mathbf{a})$ that balances accuracy, cost, convergence, and penalties, clipped to $r_{\text{max}},r_{\text{max}}]$. The Q-tables are updated as $\text{clip}(Q_{i}(s_{d},a_{i})+\alpha(\text{clip}(r(s,\mathbf{a}),-r_{\text{max}},r_{\text{max}})+\gamma\max_{a^{\prime}\in\mathcal{A}_{i}}Q_{i}(s_{d}^{\prime},a^{\prime})-Q_{i}(s_{d},a_{i})),-Q_{\text{max}},Q_{\text{max}})$, iterating over $T$ episodes and $N$ instances with decaying temperature $\tau\leftarrow\max(\tau_{\text{min}},\tau\cdot\gamma_{\tau})$.
In the inference phase, for a new instance $p$, the context $s$ is computed and discretized to $s_{d}$, and precisions are chosen greedily: $a_{i}=\arg\max_{a\in\mathcal{A}_{i}}Q_{i}(s_{d},a)$. The use of separate Q-tables reduces complexity, while clipping ensures numerical stability for ill-conditioned systems. The softmax policy and feature-based context enable the framework to adaptively optimize precision, leveraging matrix properties to achieve efficient and accurate solutions, as validated in numerical experiments in Section 5.
@@
3.3 Action Space Reduction
@@
This can naturally lead to valid precision combinations for some algorithms that is meaningful, as discussed in vi22 . This idea is similar to modified delta-debugging in 6877460 , which uses a greedy search approach instead of seeking a global minimum; it might miss potential reduced-precision configurations that suit the algorithm. The principle also naturally enables early stopping. For example, if precision $u^{\prime}_{i}$ fails, then it is not necessary to try a lower precision in step $i$ or to try lower precision for steps $j>i$ if it is already employed. Another benefit is that it allows users to define the precision selection for the $k$ computational steps in terms of their importance (the most important one is assigned to the last step $k$). One benefit of using this strategy is to reduce the action space to
$|\mathcal{A}_{\text{reduced}}|=\binom{m+k-1}{k}=\frac{m!}{k!(m-k)!}.$ (9)

### v4

# Precision autotuning for linear solvers via contextual bandit-based RL

@@ Abstract
We propose a reinforcement learning (RL) framework for adaptive precision tuning for linear solvers, which can be extended to general algorithms. The framework is formulated as a contextual bandit problem and solved using incremental action-value estimation with a discretized state space to select optimal precision configurations for computational steps, balancing precision and computational efficiency. To verify its effectiveness, we apply the framework to iterative refinement for solving linear systems $Ax=b$. In this application, our approach dynamically chooses precisions based on calculated features from the system while maintaining acceptable accuracy and convergence. In detail, an action-value estimator takes discretized features (e.g., approximate condition number and matrix norm) as input and outputs estimated action values, from which a policy selects the actions (chosen precision configurations for specific steps), optimized via an $\epsilon$-greedy strategy to maximize a multi-objective reward to balance accuracy and computational cost. Empirical results demonstrate effective precision selection, reducing computational cost while maintaining accuracy comparable to double-precision baselines. The framework generalizes to diverse out-of-sample data and provides insights into applying RL precision selection to other numerical algorithms, advancing mixed-precision numerical methods in scientific computing. To the best of our knowledge, this is the first work on precision autotuning with RL with verification on unseen datasets.
Keywords mixed-precision computing, reinforcement learning, precision tuning, linear solvers, AI for numerical methods
@@ 1 Introduction
Floating-point arithmetic (IEEE, 2019) is widely used in machine learning and scientific computing. With the rapid development of large language models (Brown et al., 2020; Hu et al., 2022; Touvron et al., 2023) as well as large-scale GPU computing (Micikevicius et al., 2017; Peng et al., 2023), numerical calculations and algorithm deployment with extensive use of floating-point arithmetic are often accompanied by a substantial increase in energy use. Therefore, using reduced-precision arithmetic in algorithms is typically required in many intensive computing kernels due to its lower communication, faster arithmetic, and energy efficiency. Seven commonly-used floating-point formats are referenced in Table 1. The NVIDIA A100 achieves up to 312 TFLOPS with half precision (FP16/BF16), 156 TFLOPS with TF32, and 9.7 TFLOPS with FP64, yielding an approximate $32\times$ speedup for low-precision formats. Further, the NVIDIA H100 (NVIDIA’s Hopper architecture) achieves up to 624 TFLOPS with fp8, offering an approximate $64\times$ speedup versus FP64 (NVIDIA Corporation, 2020; Micikevicius et al., 2022). The work Micikevicius et al. (2022) demonstrates that the FP8 formats E4M3 and E5M2 can in some cases match FP16 accuracy with reduced compute and memory costs.
@@
However, reduced-precision arithmetic often brings increased rounding errors and numerical instability, which can compromise the results of computations, especially at large scale (Lam and Hollingsworth, 2018). The idea behind mixed-precision algorithms is to design an algorithm that combines low and high precision arithmetic while retaining an acceptable level of accuracy, which remains a major obstacle in modern approximate computing paradigms (Ben Khalifa and Martel, 2023; Xu et al., 2016). Accordingly, many theoretical analyses of rounding error, mixed-precision algorithms (Carson and Higham, 2017, 2018; Carson et al., 2020; Carson and Higham, 2017; Amestoy et al., 2024), and precision tuning tools (Rubio-González et al., 2013; Guo and Rubio-González, 2018) have been developed, which aim to achieve a balance between precision loss and arithmetic speedup.
@@
Theoretical analyses of the use of multiprecision arithmetic and mixed-precision algorithms have been extensively conducted in the context of iterative refinement (Carson et al., 2020; Amestoy et al., 2024), least squares problems (Higham and Pranesh, 2021; Carson et al., 2020), and GMRES (Jang et al., 2025). For example, iterative refinement can be executed in three precisions, but the convergence is only theoretically guaranteed for a limited range of problems, which can be overly restrictive in practice Oktay and Carson (2022). As such, one can instead use to precision tuning tools to find a suitable mixed-precision approach. Existing tools typically require running the program multiple times via trial-and-error, and the scalability and general ability are not verified, so they are either limited to small or medium-scale programs or suffer from efficiency issues. An ideal case for a precision tuning tool or method is to fit the model in a limited number of training sets and learn the key features that can be used to infer mixed-precision configurations for new unseen data. These challenges need to be addressed to reduce the dependency on programmer efforts.
@@
A contextual bandit (also known as a multi-armed bandit with context) is a simplified reinforcement learning (RL) problem where the agent makes decisions based on the current context (i.e., state) without considering the future contexts or rewards of its actions; in other words, the agent optimizes for one-step (immediate) reward rather than long-term cumulative reward. In this paper, we propose a general framework for applying RL for precision selection in general numerical algorithms, formulated as a contextual bandit problem due to the independence of problem instances and the single decision per instance. This is the first work to formally define the use of RL for precision selection in any algorithm. The framework uses an epsilon-greedy strategy, controlled by a decaying epsilon determined by training episodes, to select precision levels for the prespecified algorithm steps, treating precision choices as a multi-armed bandit problem within a RL environment. The epsilon parameter balances exploration and exploitation, starting high to encourage diverse precision selections and decaying over time to favor optimal choices. We believe this framework enables at least three advantages:
@@
    Being able to inference reduced- and mixed-precision configurations for unseen datasets.
We apply this framework to the GMRES-based iterative refinement (GMRES-IR) method Carson and Higham (2017) to verify the effectiveness and generalization ability of RL for precision selection, aiming to achieve a trade-off between the use of low precision and accuracy. The framework is particularly suited for scenarios with limited training data and is designed to generalize to new data, and can be easily implemented in an online learning routine to avoid model retraining.
The rest of the paper is structured as follows. Section 2 discusses the existing literature on precision autotuning as well as its pros and cons, and a comparison with our work. Section 3 presents our general framework for precision selection; to verify its effectiveness, we use the GMRES-IR method, a well-studied mixed-precision method, as a case study detailed in Section 4. Section 5 presents the experiments and empirical results on both dense and sparse matrices under varying condition numbers and sparsity. Section 7 discusses limitations as well as future work.
@@ 2 Related Work
There exist a variety of theoretical rounding error analysis of basic mixed-precision algorithms aimed at understanding the underlying arithmetic behavior; see the survey Abdelfattah et al. (2021) and references therein. However, for real-world problems, which can be more complex, the scientific community typically uses more general approaches such as precision tuning tools, which operate at program compile time and runtime, immediately resulting in reduced-precision programs produced automatically. In this section, we briefly review existing approaches for the analysis of mixed-precision algorithms and automated precision tuning that are closely related to our method.
@@
Theoretical analysis aims to design reduced-precision settings for specific steps and analyze their rounding error so that the reduced-precision arithmetic can be used safely under certain assumptions to retain algorithm accuracy and numerical stability. Analyses in Wilkinson et al. (1966) and Moler (1967) demonstrate that in standard iterative refinement, residuals computed in extra precision (typically twice the working precision) can guarantee the convergence of the relative error to the level of the working precision if the matrix $A$ has a condition number safely less than the reciprocal of the unit-roundoff, $u$. More recently, Carson and Higham (2017) found that it is possible to produce solutions to iterative refinement with normwise relative error of order $u$ for systems with condition numbers of order $u^{-1}$ or larger by using a GMRES-based approach called GMRES-IR (see Algorithm 2). The authors extended this work in Carson and Higham (2018) to develop a general analysis of iterative refinement using up to three difference precisions, as well as a general solver. This work shows that, depending on constraints on the condition number $\kappa(A)$, the most expensive part of the computation can often be carried out entirely in low precision without harming the accuracy. However, to obtain theoretical results such as these, a great amount of effort is needed, and the analysis often relies on knowledge about the problem, e.g., the condition number, which may not be readily available in practice. To circumvent these obstacles, one can switch to precision tuning tools that operate during compile time and runtime; see the survey Cherubin and Agosta (2020) and references therein.
@@
Another approach is to use an automated precision tuning method. Such methods, built upon the software level via static and dynamic analysis tools (Ben Khalifa and Martel, 2023), can obtain suitable reduced-precision settings through trial and error. The search-based tool Precimonious (Rubio-González et al., 2013) uses high-performance dynamic programming approaches, which achieve a potentially lower precision setting for the program via Delta Debugging of the space of possible floating-point configurations within the target accuracy. Further, HiFPTuner (Guo and Rubio-González, 2018) (High-level Floating-Point Tuner) operates at the source level to automatically tune the floating-point precision via profiling floating-point operations and estimating the numerical impact of the reduced-precision program. CRAFT (Lam and Hollingsworth, 2018) builds reduced-precision configurations of floating-point programs based on the Dyninst binary instrumentation framework (Buck and Hollingsworth, 2000), and uses an automatic search process that identifies the precision-level sensitivity of various parts of an application. Search-based tools for finding a low-precision configuration typically require a dramatic increase in computational cost as the number of variables increases, and do not scale well to large programs. To further reduce the runtime of precision tuning, Blame Analysis (Rubio-González et al., 2016) is devoted to executing a program once, analyzing floating-point instructions to determine the minimum precision required for each operand, reducing the search space of Precimonious and improving scalability. PyFloT (Brunie et al., 2020) automates mixed-precision tuning for HPC applications by leveraging runtime profiling, instruction-level granularity, and temporal locality to learn numerical behavior and recurring patterns in instruction execution. FPLearner (Wang and Rubio-González, 2024) leverages a graph representation learning technique to predict the performance and accuracy of a mixed-precision program, which can then be used to accelerate dynamic precision tuning by reducing the number of program runs.
@@ 3.1 Contextual Bandit Formulation
3.2 Learning with Discretized Context Space
@@
$s_{d}=\text{discretize}(s)=\left(\text{bin}(\phi_{1}(p)),\ldots,\text{bin}(\phi_{d}(p)))\right),$ (4)
@@
To handle the combinatorial action space $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{k}$ with cardinality $|\mathcal{A}|=m^{k}$, we directly learn a tabular action-value function
$Q:\mathcal{S}_{d}\times\mathcal{A}\to\mathbb{R},$
where $Q(s_{d},\mathbf{a})$ estimates the expected reward obtained by applying the precision configuration $\mathbf{a}=(a_{1},\ldots,a_{k})\in\mathcal{A}$ to the solver under discretized state $s_{d}\in\mathcal{S}_{d}$.
This formulation treats each precision configuration as a distinct action, capturing the joint effect of precision choices across all computational steps. The resulting Q-table has size $|\mathcal{S}_{d}|\cdot|\mathcal{A}|=|\mathcal{S}_{d}|\cdot m^{k}$, which can become prohibitively large as the number of steps $k$ or precision levels $m$ increases.
In the following subsection, we introduce a structured reduction of the action space to mitigate this exponential growth while preserving numerically meaningful configurations.
@@
Algorithm 1 Contextual Bandit Framework for Precision Selection
1:Input: Training instances $\mathcal{P}=\{p_{1},\ldots,p_{N}\}$, numerical method $\mathcal{M}$, precision sets $\{\mathcal{A}_{i}\}_{i=1}^{k}$, reward function $r$, episodes $T$, learning-rate schedule $\alpha_{t}$, exploration rate $\epsilon$, bin sets $\{\Phi_{j}\}_{j=1}^{d}$
2:Construct the joint action space $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{k}$
3:Initialize a tabular action-value estimator $Q(s_{d},\mathbf{a})$ for all $s_{d}\in\mathcal{S}_{d}$ and $\mathbf{a}\in\mathcal{A}$
4:Initialize visit counts $N(s_{d},\mathbf{a})\leftarrow 0$ for all $s_{d}\in\mathcal{S}_{d}$ and $\mathbf{a}\in\mathcal{A}$
5:for episode $t=1$ to $T$ do
6:  for each instance $p_{i}\in\mathcal{P}$ do
7:   Compute context $s_{i}\leftarrow[\phi_{1}(p_{i}),\ldots,\phi_{d}(p_{i})]$
8:   Discretize context $s_{d}\leftarrow\mathrm{discretize}(s_{i},\{\Phi_{j}\})$
9:   Select action $\mathbf{a}\in\mathcal{A}$ using an $\epsilon$-greedy policy:
@@
10:   Run $\mathcal{M}$ with precision configuration $\mathbf{a}$ on $p_{i}$
11:   Observe reward $r_{i}\leftarrow r(s_{d},\mathbf{a})$
12:   Update visit count: $N(s_{d},\mathbf{a})\leftarrow N(s_{d},\mathbf{a})+1$
13:   Set learning rate, e.g.,
$\alpha\leftarrow\frac{1}{N(s_{d},\mathbf{a})}$
14:   Update the tabular action-value estimator:
$Q(s_{d},\mathbf{a})\leftarrow Q(s_{d},\mathbf{a})+\alpha\big(r_{i}-Q(s_{d},\mathbf{a})\big)$
15:  end for
16:  Update $\epsilon$
17:end for
18:Inference: For a new instance $p$, compute $s=[\phi_{1}(p),\ldots,\phi_{d}(p)]$, discretize $s_{d}=\mathrm{discretize}(s,\{\Phi_{j}\})$, and select
$\mathbf{a}^{*}=\arg\max_{\mathbf{a}\in\mathcal{A}}Q(s_{d},\mathbf{a}).$
19:Apply $\mathcal{M}$ with precision configuration $\mathbf{a}^{*}$
Our learning framework for precision selection is formulated as a contextual bandit over a discretized context space. For each problem instance $p=(A,b)\in\mathcal{P}$, we compute a feature vector $s=[\phi_{1}(p),\ldots,\phi_{d}(p)]$ and map it to a discrete context $s_{d}\in\mathcal{S}_{d}$. The agent then selects a joint precision configuration $\mathbf{a}=(a_{1},\ldots,a_{k})\in\mathcal{A}$ for the $k$ computational steps of the solver. After executing the mixed-precision algorithm with action $\mathbf{a}$, the environment returns a scalar reward $r(s_{d},\mathbf{a})$ that balances numerical accuracy, computational cost, convergence behavior, and iteration count.
To estimate the expected reward of each state-action pair, we maintain a single tabular function $Q(s_{d},\mathbf{a})$. During training, actions are sampled using an $\epsilon$-greedy exploration rule, and the table is updated using the incremental estimator
$Q(s_{d},\mathbf{a})\leftarrow Q(s_{d},\mathbf{a})+\alpha_{t}(s_{d},\mathbf{a})\big(r(s_{d},\mathbf{a})-Q(s_{d},\mathbf{a})\big).$
In the inference phase, for a new instance $p$, the context $s$ is computed and discretized to $s_{d}$, and the precision configuration is selected greedily as
$\mathbf{a}^{*}(s_{d})=\arg\max_{\mathbf{a}\in\mathcal{A}}Q(s_{d},\mathbf{a}).$
Using a single Q-table over the reduced joint action space preserves the interaction among precision choices across different computational steps while keeping the learning problem tractable. The feature-based context enables the framework to adaptively select numerically effective and computationally efficient precision configurations, as validated in the numerical experiments in Section 5.
@@
Theoretical Justification of Discretization
Our contextual bandit framework relies on the discretization of the continuous context space $\mathcal{S}$ into a finite set $\mathcal{S}_{d}$. This raises the question of how much optimality is lost due to this approximation. We provide a bound below showing that, under a mild smoothness assumption on the expected reward, the performance loss induced by discretization is controlled by the granularity of the bins.
Assumption 1 (Lipschitz continuity of the expected reward).
Let $\mu(s,\mathbf{a}):=\mathbb{E}[r(s,\mathbf{a})]$ denote the expected reward. We assume, for all $\mathbf{a}\in\mathcal{A}$, there exists a constant $L>0$ such that
$|\mu(s,\mathbf{a})-\mu(s^{\prime},\mathbf{a})|\leq L\|s-s^{\prime}\|,\quad\forall s,s^{\prime}\in\mathcal{S}.$ (8)
Proposition 1 (Discretization error bound).
Let $s\in\mathcal{S}$ be a continuous context, and let $s_{d}\in\mathcal{S}_{d}$ be its discretized representation. Assume that each discretization bin has diameter at most $\Delta$, i.e.,
$\|s-\omega(s_{d})\|\leq\Delta,$
where $\omega(s_{d})$ denotes a representative point (e.g., bin center) of $s_{d}$.
Define the optimal action in the continuous space and the discretized policy, respectively, as
$\mathbf{a}^{*}(s)=\arg\max_{\mathbf{a}\in\mathcal{A}}\mu(s,\mathbf{a}),\qquad\mathbf{a}_{d}^{*}(s_{d})=\arg\max_{\mathbf{a}\in\mathcal{A}}\mu(\omega(s_{d}),\mathbf{a}).$
Under Assumption 1, the performance loss due to discretization is bounded by
$\mu(s,\mathbf{a}^{*}(s))-\mu(s,\mathbf{a}_{d}^{*}(s_{d}))\leq 2L\Delta.$ (9)
Proof.
Let $\omega=\omega(s_{d})$ be the representative point associated with the discretized state $s_{d}$. By Assumption 1, for any $\mathbf{a}\in\mathcal{A}$, we have
$|\mu(s,\mathbf{a})-\mu(\omega,\mathbf{a})|\leq L\|s-\omega\|\leq L\Delta.$ (10)
By definition of $\mathbf{a}^{*}(s)$,
$\mu(s,\mathbf{a}^{*}(s))\leq\mu(\omega,\mathbf{a}^{*}(s))+L\Delta.$
Since $\mathbf{a}_{d}^{*}(s_{d})$ maximizes $\mu(\omega,\mathbf{a})$ over $\mathcal{A}$, we have
$\mu(\omega,\mathbf{a}^{*}(s))\leq\mu(\omega,\mathbf{a}_{d}^{*}(s_{d})).$
Applying the Lipschitz bound again,
$\mu(\omega,\mathbf{a}_{d}^{*}(s_{d}))\leq\mu(s,\mathbf{a}_{d}^{*}(s_{d}))+L\Delta.$
Combining the above inequalities yields
$\mu(s,\mathbf{a}^{*}(s))\leq\mu(s,\mathbf{a}_{d}^{*}(s_{d}))+2L\Delta,$
which proves the result. ∎
Proposition 1 shows that the loss induced by discretizing the context space is controlled by the bin diameter $\Delta$. In particular, if the expected reward varies smoothly with respect to problem features (such as condition number and matrix norm), then finer discretizations lead to policies that approach the optimal continuous-context policy. This provides a theoretical justification for the discretization strategy used in our contextual bandit framework.
Action Space Reduction
@@
This can naturally lead to valid precision combinations for some algorithms that is meaningful, as discussed in Vieublé (2022). This idea is similar to modified delta-debugging in Rubio-González et al. (2013), which uses a greedy search approach instead of seeking a global minimum; it might miss potential reduced-precision configurations that suit the algorithm. The principle also naturally enables early stopping. For example, if precision $u^{\prime}_{i}$ fails, then it is not necessary to try a lower precision in step $i$ or to try lower precision for steps $j>i$ if it is already employed. Another benefit is that it allows users to define the precision selection for the $k$ computational steps in terms of their importance (the most important one is assigned to the last step $k$). One benefit of using this strategy is to reduce the action space to
$|\mathcal{A}_{\text{reduced}}|=\binom{m+k-1}{k}=\frac{(m+k-1)!}{k!(m-1)!}.$ (12)